Please I need help with these 5 problems. Thank you.

1. Write an algebraic expression for the following. Double a number plus 7

6 / 1
- - = (6 over 8 divided by 1 over 4)
8 4

5 - 3
- = (5 - 3 over 8)
8

7 3 - 3 1
- - = ( 7 and 3 over 4 (-minus) 3 and 1 over 2)
4 2

4 x 4
- - = (4 over 11 multiply by 4 over 6)
11 6

Answer by stanbon(75887)

1. Write an algebraic expression for the following. Double a number plus 7
2x + 7
--------------
6 / 1
- - = (6 over 8 divided by 1 over 4)
8 4
---
(6/8)/(1/4)
= (3/4)*(4/1)
= 3
-------------
5 - 3
- = (5 - 3 over 8)
8
---
(5/8) - 3
(5/8) - (24/8)
(-19/8)
------------------
7 3 - 3 1
- - = ( 7 and 3 over 4 (-minus) 3 and 1 over 2)
4 2
---
(7 + (3/4)) - (3 + (1/2))
(28/4 + 3/4) - (6/2 + 1/2)
(31/4) - (7/2)
(31/4) - (14/3)
= 17/4
===================
4 x 4
- - = (4 over 11 multiply by 4 over 6)
11 6
---
(4/11) * (4/6)
(4/11) * (2/3)
8/33
=======================
Cheers,
Stan H.